OpenStudy (anonymous):

One leg of a right triangle measures 96 inches. The length of the hypotenuse is twice the other leg. If x represents the length of the other leg, which of the following equations is the Pythagorean Theorem written in general form? -3x^2 - 9216 = 0 4x^2 - 9216 = 0 3x^2 - 9216 = 0

OpenStudy (anonymous):

...

563blackghost (563blackghost):

the pythagorean formula is a^2 + b^2 = c^2 a and b are legs and c is the hypotenuse. let a = 96 let b = x let c = 2x by the pythagorean formula, 96^2 + x^2 = (2x)^2 which becomes 96^2 + x^2 = 4x^2 subtract x^2 from both sides of the equation to get 96^2 = 4x^2 - x^2 which becomes 96^2 = 3x^2 take the square root of both sides of the equation to get 96 = x*sqrt(3) that makes x = 96/sqrt(3).

OpenStudy (anonymous):

I already found that same answer on some other website..lol still didn't help.

OpenStudy (anonymous):

Is it B? @563blackghost

OpenStudy (loser66):



563blackghost (563blackghost):

96/sqrt(3) is 55.42562584 if you input this decimal into the answer equations one will equal 0.....so A is -18432.....B is 3072 .... and C is 0...